# Shallow Flow Matching for Coarse-to-Fine Text-to-Speech Synthesis

### v1

@@ 5.2 Acceleration of adaptive-step ODE solvers
Table 5: Partial RTF results for adaptive-step ODE solvers. Complete results are provided in Appendix F. $\overline{\text{RTF}}$ denotes the mean RTF. Rate (%) denotes the relative speedup compared to the ablated model.
@@
Ablated 0.401 0.000 0.056 0.000 0.270 0.000 0.145 0.000
SFM ($\alpha$=1.0) 0.361 9.858 0.053 4.171 0.225 16.924 0.132 9.100
SFM ($\alpha$=2.0) 0.299 25.541 0.048 13.306 0.187 30.931 0.115 20.484
SFM ($\alpha$=3.0) 0.249 37.931 0.043 22.749 0.157 41.895 0.100 30.753
SFM ($\alpha$=4.0) 0.207 48.486 0.037 34.246 0.131 51.576 0.086 40.559
SFM ($\alpha$=5.0) 0.157 60.825 0.027 50.968 0.110 59.266 0.076 47.605
In Table 5, we only show the mean RTF for clarity in the five run times. We also show the speedup rate compared to the ablated model. From Table 5 and the tables in Appendix F, we observe that increasing $\alpha$ improves the signal-to-noise ratio of the initial state during flow inference, which stabilizes the ODE solving process. As a result, fewer forward steps are required, and the overall inference time is significantly reduced. Notably, this acceleration effect is limited to adaptive-step solvers, as fixed-step solvers perform a predefined number of steps, and thus cannot leverage the improved stability of the initial stages to reduce inference cost.
@@ 7 Limitations
This work applies only minimal modifications to the backbone TTS models, using a simple SFM head to demonstrate the effectiveness of our proposed SFM method. Therefore, there is room for improvement and extension of the SFM framework. For example, a more powerful SFM head could be used when the weak generator is unstable. In addition, the flow conditions could be enhanced by incorporating text or speech token embeddings to better align the final outputs with the input semantic features.
Although our method is designed for the coarse-to-fine TTS models, it also holds potential in other domains, such as speech denoising and enhancement, or image generation tasks like denoising and super-resolution. Its effectiveness in other domains still needs to be validated.
@@ 8 Conclusion
We introduce a novel Shallow Flow Matching (SFM) method for coarse-to-fine TTS models. SFM leverages coarse representations to construct intermediate states on the CondOT path, from which flow-based inference produces more stable generation, more natural synthetic speech, and faster inference when using adaptive-step ODE solvers. The proposed method also has the potential to be extended to other tasks. We hope it inspires future explorations of correcting or enhancing the early stages of flow matching generation.
@@ Appendix A Theorem Proofs
$\boldsymbol{x}_{\tau}$ =$=\begin{cases}\sqrt{(1-(1-\sigma_{\mathrm{min}})t_{m})^{2}-\sigma^{2}_{m}}\boldsymbol{x}_{0}+\boldsymbol{x}_{m},&\text{if }\Delta<1,\\ \frac{1}{\Delta}\boldsymbol{x}_{m},&\text{if }\Delta\geq 1,\end{cases}$ (24)
@@
$\tau$ =$=\begin{cases}t_{m},&\text{if }\Delta<1,\\ \frac{1}{\Delta}t_{m},&\text{if }\Delta\geq 1.\end{cases}$ (25)
$\boldsymbol{\mu}(\boldsymbol{x}_{\tau})$ =$=\begin{cases}t_{m}\boldsymbol{x}_{1}=\tau\boldsymbol{x}_{1},&\text{if }\Delta<1,\\ \frac{1}{\Delta}t_{m}\boldsymbol{x}_{1}=\tau\boldsymbol{x}_{1},&\text{if }\Delta\geq 1,\end{cases}$ (26)
$\sigma(\boldsymbol{x}_{\tau})$ =$=\begin{cases}\sqrt{(1-(1-\sigma_{\mathrm{min}})t_{m})^{2}-\sigma^{2}_{m}+\sigma^{2}_{m}}=1-(1-\sigma_{\mathrm{min}})\tau,&\text{if }\Delta<1,\\ \frac{\sigma_{m}}{\Delta}=1-(1-\sigma_{\mathrm{min}})\frac{t_{m}}{\Delta}=1-(1-\sigma_{\mathrm{min}})\tau,&\text{if }\Delta\geq 1,\end{cases}$ (27)
@@ III. When $\sigma_{m}>1-(1-\sigma_{\mathrm{min}})t_{m}$, then $\Delta>1$:
Thus, $T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m})$ is continuous in the Wasserstein-2 metric. ∎
@@
$\boldsymbol{x}_{t}$ =$=\begin{cases}(1-\frac{t}{t_{m}})\boldsymbol{x}_{0}+\frac{t}{t_{m}}\boldsymbol{x}_{t_{m}},&\text{if }t<t_{m},\\ (1-\frac{t-t_{m}}{1-t_{m}})\boldsymbol{x}_{t_{m}}+\frac{t-t_{m}}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}),&\text{if }t\geq t_{m},\end{cases}$ (62)
$\boldsymbol{u}_{t}$ =$=\begin{cases}\frac{1}{t_{m}}(\boldsymbol{x}_{t_{m}}-\boldsymbol{x}_{0}),&\text{if }t<t_{m},\\ \frac{1}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}-\boldsymbol{x}_{t_{m}}),&\text{if }t\geq t_{m}.\end{cases}$ (63)
@@ Appendix C Algorithms
The SFM-head $\boldsymbol{h_{\psi}}$ predicts $\log\hat{\sigma}^{2}_{\boldsymbol{h}}$ instead of $\hat{\sigma}^{2}_{\boldsymbol{h}}$ directly, in order to ensure numerical stability and positive outputs.
We also modify the placement of the $\mathcal{L}_{\mu}$ for implementation convenience. Compared to the formulation in Equation 13, this modification results in a scaled version of the original loss by a constant factor of $\frac{1}{\Delta^{2}}$.
@@
$\mathcal{L}_{\mathrm{coarse}}=\mathbb{E}||\boldsymbol{X}_{\boldsymbol{g}}-\boldsymbol{X}_{1}||^{2}$
@@
$t_{\boldsymbol{h}}\leftarrow\mathbb{E}\left[\frac{sg[\boldsymbol{X}_{\boldsymbol{h}}]\cdot\boldsymbol{X}_{1}}{\boldsymbol{X}_{1}\cdot\boldsymbol{X}_{1}}\right],\quad\sigma_{\boldsymbol{h}}^{2}\leftarrow\mathbb{E}||sg[{\boldsymbol{X}}_{\boldsymbol{h}}]-t_{\boldsymbol{h}}\boldsymbol{X}_{1}||^{2}$
@@
$\mathcal{L}_{t}=(\hat{t}_{\boldsymbol{h}}-t_{\boldsymbol{h}})^{2},\quad\mathcal{L}_{\sigma}=(\log\hat{\sigma}^{2}_{\boldsymbol{h}}-\log\sigma^{2}_{\boldsymbol{h}})^{2},\quad\mathcal{L}_{\mu}\ =\mathbb{E}||\boldsymbol{X}_{\boldsymbol{h}}-t_{\boldsymbol{h}}\boldsymbol{X}_{1}||^{2}$
@@
$\mathcal{L}_{\mathrm{CFM}}=\mathbb{E}||\boldsymbol{v_{\theta}}(\boldsymbol{X}_{t},t)-\boldsymbol{U}_{t}||^{2}$
@@
[…]
@@ Appendix D Complete results of $\boldsymbol{\alpha}$ selection
We provide complete results in optimal $\alpha$ selection for all models in this section, including Table 6, Table 7, Table 8, and Table 9. Notably, $\hat{t}_{\boldsymbol{g}}$ and $\hat{\sigma}_{\boldsymbol{g}}$ denote the mean values computed over all utterances in the validation sets.
@@ Appendix E Minor analysis of evaluation results
Appendix F RTF with using adaptive-step ODE solvers
@@
Table 10: RTF results for adaptive-step ODE solvers (solver inference). RTF ($\mu\pm\sigma$) denotes the mean and std of RTF. Rate (%) denotes the relative speedup compared to the ablated model.
@@
Ablated 0.401±0.001 0.000 0.056±0.000 0.000 0.270±0.001 0.000 0.145±0.000 0.000
SFM ($\alpha$=1.0) 0.361±0.001 9.858 0.053±0.000 4.171 0.225±0.001 16.924 0.132±0.000 9.100
SFM ($\alpha$=2.0) 0.299±0.001 25.541 0.048±0.000 13.306 0.187±0.001 30.931 0.115±0.001 20.484
SFM ($\alpha$=3.0) 0.249±0.002 37.931 0.043±0.000 22.749 0.157±0.001 41.895 0.100±0.000 30.753
SFM ($\alpha$=4.0) 0.207±0.001 48.486 0.037±0.000 34.246 0.131±0.000 51.576 0.086±0.001 40.559
SFM ($\alpha$=5.0) 0.157±0.000 60.825 0.027±0.000 50.968 0.110±0.000 59.266 0.076±0.000 47.605
@@
Ablated 0.952±0.002 0.000 0.131±0.001 0.000 0.690±0.001 0.000 0.371±0.001 0.000
SFM ($\alpha$=1.0) 0.950±0.001 0.210 0.123±0.000 6.269 0.690±0.003 0.030 0.371±0.001 0.111
SFM ($\alpha$=2.0) 0.771±0.002 18.998 0.111±0.001 15.463 0.558±0.001 19.071 0.318±0.001 14.356
SFM ($\alpha$=3.0) 0.626±0.001 34.271 0.100±0.001 23.635 0.436±0.001 36.844 0.265±0.001 28.626
SFM ($\alpha$=4.0) 0.519±0.001 45.449 0.093±0.000 29.503 0.350±0.001 49.257 0.231±0.000 37.846
SFM ($\alpha$=5.0) 0.407±0.002 57.242 0.073±0.001 44.413 0.279±0.001 59.547 0.195±0.001 47.454
@@
Ablated 2.153±0.007 0.000 0.394±0.002 0.000 1.270±0.005 0.000 0.998±0.002 0.000
SFM ($\alpha$=1.0) 1.212±0.001 43.694 0.261±0.000 33.707 0.790±0.001 37.823 0.690±0.002 30.849
SFM ($\alpha$=2.0) 0.982±0.002 54.386 0.232±0.000 41.301 0.697±0.001 45.123 0.613±0.001 38.593
SFM ($\alpha$=3.0) 0.789±0.000 63.346 0.205±0.000 48.081 0.611±0.001 51.930 0.564±0.000 43.473
SFM ($\alpha$=4.0) 0.592±0.001 72.497 0.169±0.000 57.149 0.542±0.001 57.305 0.489±0.000 51.003
SFM ($\alpha$=5.0) 0.586±0.001 72.792 0.173±0.000 56.096 0.538±0.000 57.674 0.505±0.001 49.377
@@
Ablated 2.532±0.012 0.000 0.369±0.001 0.000 2.175±0.006 0.000 1.397±0.005 0.000
SFM ($\alpha$=1.0) 1.745±0.004 31.093 0.283±0.001 23.449 1.454±0.002 33.143 1.061±0.013 24.069
SFM ($\alpha$=2.0) 1.426±0.006 43.667 0.250±0.000 32.196 1.211±0.006 44.326 0.917±0.002 34.395
SFM ($\alpha$=3.0) 1.110±0.004 56.182 0.219±0.000 40.682 1.001±0.003 53.952 0.787±0.002 43.679
SFM ($\alpha$=4.0) 0.963±0.001 61.967 0.186±0.000 49.642 0.887±0.001 59.217 0.720±0.002 48.469
SFM ($\alpha$=5.0) 0.960±0.002 62.096 0.185±0.000 49.875 0.888±0.001 59.190 0.727±0.002 47.990
@@
Ablated 1.170±0.002 0.000 0.188±0.000 0.000 0.699±0.002 0.000 0.539±0.003 0.000
SFM ($\alpha$=1.0) 0.621±0.001 46.934 0.125±0.000 33.704 0.396±0.003 43.324 0.337±0.001 37.591
SFM ($\alpha$=2.0) 0.503±0.001 56.980 0.108±0.001 42.475 0.336±0.000 51.894 0.286±0.001 46.910
SFM ($\alpha$=3.0) 0.391±0.001 66.558 0.092±0.000 51.292 0.295±0.000 57.811 0.257±0.000 52.252
SFM ($\alpha$=4.0) 0.300±0.000 74.320 0.073±0.000 61.121 0.259±0.000 62.958 0.229±0.000 57.595
SFM ($\alpha$=5.0) 0.302±0.000 74.212 0.074±0.000 60.825 0.258±0.001 63.061 0.229±0.000 57.523
Table 11: RTF results when using adaptive-step ODE solvers (model inference). RTF ($\mu\pm\sigma$) denotes the mean and std of RTF. Rate (%) denotes the relative speedup compared to the ablated model.
@@
Ablated 0.403±0.001 0.000 0.057±0.000 0.000 0.272±0.001 0.000 0.146±0.000 0.000
SFM ($\alpha$=1.0) 0.363±0.001 9.809 0.055±0.000 4.017 0.226±0.001 16.814 0.133±0.000 8.976
SFM ($\alpha$=2.0) 0.300±0.001 25.430 0.050±0.000 12.866 0.188±0.001 30.735 0.117±0.001 20.230
SFM ($\alpha$=3.0) 0.251±0.002 37.773 0.045±0.000 22.063 0.159±0.001 41.639 0.102±0.000 30.399
SFM ($\alpha$=4.0) 0.208±0.001 48.282 0.038±0.000 33.261 0.133±0.000 51.268 0.088±0.001 40.104
SFM ($\alpha$=5.0) 0.159±0.000 60.578 0.029±0.000 49.541 0.112±0.000 58.914 0.077±0.000 47.076
@@
Ablated 0.955±0.002 0.000 0.135±0.001 0.000 0.694±0.001 0.000 0.375±0.001 0.000
SFM ($\alpha$=1.0) 0.953±0.001 0.211 0.127±0.000 6.111 0.693±0.003 0.036 0.374±0.001 0.116
SFM ($\alpha$=2.0) 0.775±0.002 18.925 0.115±0.001 15.046 0.562±0.001 18.971 0.321±0.001 14.213
SFM ($\alpha$=3.0) 0.629±0.001 34.141 0.104±0.001 23.000 0.439±0.001 36.651 0.268±0.001 28.349
SFM ($\alpha$=4.0) 0.523±0.001 45.274 0.096±0.000 28.711 0.354±0.001 48.998 0.234±0.000 37.481
SFM ($\alpha$=5.0) 0.411±0.002 57.021 0.077±0.001 43.209 0.283±0.001 59.235 0.199±0.001 46.993
@@
Ablated 2.157±0.007 0.000 0.398±0.002 0.000 1.274±0.005 0.000 1.001±0.002 0.000
SFM ($\alpha$=1.0) 1.216±0.001 43.622 0.265±0.000 33.396 0.793±0.001 37.716 0.694±0.002 30.730
SFM ($\alpha$=2.0) 0.986±0.002 54.294 0.235±0.000 40.912 0.701±0.001 44.992 0.616±0.001 38.447
SFM ($\alpha$=3.0) 0.793±0.000 63.242 0.208±0.000 47.643 0.614±0.001 51.781 0.568±0.000 43.315
SFM ($\alpha$=4.0) 0.596±0.001 72.380 0.172±0.000 56.640 0.546±0.001 57.146 0.492±0.000 50.821
SFM ($\alpha$=5.0) 0.589±0.001 72.674 0.177±0.000 55.598 0.541±0.000 57.513 0.509±0.001 49.200
@@
Ablated 2.533±0.012 0.000 0.370±0.001 0.000 2.176±0.006 0.000 1.398±0.005 0.000
SFM ($\alpha$=1.0) 1.746±0.004 31.076 0.284±0.001 23.353 1.455±0.002 33.124 1.062±0.013 24.035
SFM ($\alpha$=2.0) 1.428±0.006 43.644 0.252±0.000 32.070 1.212±0.006 44.301 0.918±0.002 34.360
SFM ($\alpha$=3.0) 1.111±0.004 56.155 0.220±0.000 40.536 1.003±0.003 53.922 0.788±0.002 43.638
SFM ($\alpha$=4.0) 0.964±0.001 61.937 0.187±0.000 49.471 0.888±0.001 59.184 0.721±0.002 48.424
SFM ($\alpha$=5.0) 0.961±0.002 62.066 0.186±0.000 49.705 0.889±0.001 59.157 0.728±0.002 47.946
@@
Ablated 1.171±0.002 0.000 0.190±0.000 0.000 0.700±0.002 0.000 0.541±0.003 0.000
SFM ($\alpha$=1.0) 0.622±0.001 46.870 0.126±0.000 33.406 0.397±0.003 43.221 0.338±0.001 37.472
SFM ($\alpha$=2.0) 0.505±0.001 56.903 0.110±0.001 42.113 0.338±0.000 51.779 0.288±0.001 46.775
SFM ($\alpha$=2.0) 0.393±0.001 66.472 0.093±0.000 50.872 0.296±0.000 57.683 0.259±0.000 52.102
SFM ($\alpha$=4.0) 0.302±0.000 74.225 0.075±0.000 60.633 0.260±0.000 62.820 0.230±0.000 57.432
SFM ($\alpha$=5.0) 0.303±0.000 74.116 0.075±0.000 60.336 0.260±0.001 62.924 0.230±0.000 57.360

### v2

@@ 5.2 Acceleration of adaptive-step ODE solvers
Table 5: Partial RTF and NFE results for adaptive-step ODE solvers. Complete results are provided in Appendix F. $\overline{\text{RTF}}$ denotes the mean RTF. Rate (%) denotes the relative speedup in terms of $\overline{\text{RTF}}$ compared to the ablated model.
@@
Baseline 0.407 -1.496 312.36 0.054 3.571 44.82 0.271 -0.370 223.95 0.142 2.069 120.10
Ablated 0.401 0.000 306.72 0.056 0.000 45.28 0.270 0.000 221.81 0.145 0.000 121.46
SFM ($\alpha$=1.0) 0.361 9.858 277.56 0.053 4.171 44.08 0.225 16.924 186.23 0.132 9.100 111.14
SFM ($\alpha$=2.0) 0.299 25.541 229.43 0.048 13.306 39.16 0.187 30.931 153.08 0.115 20.484 96.02
SFM ($\alpha$=3.0) 0.249 37.931 191.49 0.043 22.749 34.88 0.157 41.895 129.02 0.100 30.753 84.20
SFM ($\alpha$=4.0) 0.207 48.486 156.78 0.037 34.246 29.74 0.131 51.576 107.90 0.086 40.559 72.56
SFM ($\alpha$=5.0) 0.157 60.825 120.07 0.027 50.968 22.14 0.110 59.266 90.74 0.076 47.605 63.74
Table 5 and the tables in Appendix F report only the mean RTF for clarity, as the standard deviations across the five runs are all below 0.013. We also present the speedup rate in terms of RTF, measured relative to the ablated model. From these tables, we observe that increasing $\alpha$ improves the signal-to-noise ratio of the initial state during flow inference, which stabilizes the ODE solving process. As a result, fewer forward steps are required, and the overall inference time is significantly reduced. Notably, this acceleration effect is limited to adaptive-step solvers, as fixed-step solvers perform a predefined number of steps, and thus cannot leverage the improved stability of the initial stages to reduce inference cost.
@@ 7 Limitations
This work applies only minimal modifications to the backbone TTS models and uses a simple SFM head to demonstrate the effectiveness of our proposed SFM method. Therefore, there is considerable room for improving the SFM framework and corresponding implementation. For example:
1. A more powerful SFM head could be used when the weak generator is unstable.
2. The flow conditions could be enhanced by incorporating text or speech token embeddings to better align the final outputs with the input semantic features.
3. When applying SFM on CosyVoice, we directly concatenate the speech tokens, speaker embeddings, and masked mel-spectrograms as the encoder input. This naive fusion strategy may negatively affect cross-modal alignment and calls for a more elaborate integration approach.
@@ 8 Conclusion
We introduce a novel Shallow Flow Matching (SFM) method for coarse-to-fine TTS models. SFM leverages coarse representations to construct intermediate states on the CondOT path, enabling flow-based inference to produce more stable generation, more natural synthetic speech, and faster inference when using adaptive-step ODE solvers.
SFM focuses on cases where the FM module serves as a refiner. Although it is validated on TTS tasks, the underlying framework and theoretical foundation are general. It also holds potential for other domains, such as speech denoising and enhancement, as well as image generation tasks like denoising and super-resolution. Further exploration of its applications to these tasks remains a promising direction for future work.
9 Acknowledgements
This research was supported by JST Moonshot Grant Number JPMJMS2237 and JST SPRING Grant Number JPMJSP2108.
@@ Appendix A Theorem Proofs
$\boldsymbol{x}_{\tau}$ $=\begin{cases}\sqrt{(1-(1-\sigma_{\mathrm{min}})t_{m})^{2}-\sigma^{2}_{m}}\boldsymbol{x}_{0}+\boldsymbol{x}_{m},&\text{if }\Delta<1,\\ \frac{1}{\Delta}\boldsymbol{x}_{m},&\text{if }\Delta\geq 1,\end{cases}$ (24)
@@
[…]
$\sigma(\boldsymbol{x}_{\tau})$ $=\begin{cases}\sqrt{(1-(1-\sigma_{\mathrm{min}})t_{m})^{2}-\sigma^{2}_{m}+\sigma^{2}_{m}}=1-(1-\sigma_{\mathrm{min}})\tau,&\text{if }\Delta<1,\\ \frac{\sigma_{m}}{\Delta}=1-(1-\sigma_{\mathrm{min}})\frac{t_{m}}{\Delta}=1-(1-\sigma_{\mathrm{min}})\tau,&\text{if }\Delta\geq 1,\end{cases}$ (27)
@@ III. When $\sigma_{m}>1-(1-\sigma_{\mathrm{min}})t_{m}$, then $\Delta>1$:
Thus, $T_{\boldsymbol{x}_{1}}(t_{m},\sigma_{m})$ is continuous with respect to $t_{m}$ and $\sigma_{m}$ in the Wasserstein-2 metric. ∎
@@
$\boldsymbol{x}_{t}$ $=\begin{cases}(1-\frac{t}{t_{m}})\boldsymbol{x}_{0}+\frac{t}{t_{m}}\boldsymbol{x}_{t_{m}},&\text{if }t<t_{m},\\ (1-\frac{t-t_{m}}{1-t_{m}})\boldsymbol{x}_{t_{m}}+\frac{t-t_{m}}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}),&\text{if }t\geq t_{m},\end{cases}$ (62)
$\boldsymbol{u}_{t}$ $=\begin{cases}\frac{1}{t_{m}}(\boldsymbol{x}_{t_{m}}-\boldsymbol{x}_{0}),&\text{if }t<t_{m},\\ \frac{1}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}-\boldsymbol{x}_{t_{m}}),&\text{if }t\geq t_{m}.\end{cases}$ (63)
@@ Appendix C Algorithms
In the practical implementation, we apply several detailed modifications and tricks:
1. The SFM-head $\boldsymbol{h_{\psi}}$ predicts $\log\hat{\sigma}^{2}_{\boldsymbol{h}}$ instead of $\hat{\sigma}^{2}_{\boldsymbol{h}}$ directly to ensure numerical stability and positive outputs.
2. Due to the $\mathcal{L}_{\mathrm{coarse}}$, the $t_{\boldsymbol{h}}$ obtained from $\boldsymbol{X}_{\boldsymbol{h}}$ are generally larger than 0 during training. Therefore, we remove the constraint $t_{\boldsymbol{h}}\geq 0$ in Equation 13.
3. We also modify the placement of the $\mathcal{L}_{\mu}$ for implementation convenience. Compared to the formulation in Equation 13, this modification results in a scaled version of the original loss by a constant factor of $\frac{1}{\Delta^{2}}$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\boldsymbol{X}_{\mathrm{pred}}=\boldsymbol{X}_{\boldsymbol{h}}+\int_{t_{\boldsymbol{h}}}^{1}\boldsymbol{v_{\theta}}(\boldsymbol{X}_{t},t)\,dt$
@@ Appendix D Complete results of $\boldsymbol{\alpha}$ selection
We provide complete results in optimal $\alpha$ selection for all models in this section, including Table 6, Table 7, Table 8, and Table 9.
@@ Appendix E Minor analysis of evaluation results
Appendix F RTF and NFE with using adaptive-step ODE solvers
@@
Table 10: RTF and NFE results for adaptive-step ODE solvers (solver inference). $\overline{\text{RTF}}$ denotes the mean of RTF. Rate (%) denotes the relative speedup in terms of $\overline{\text{RTF}}$ compared to the ablated model.
@@
Baseline 0.407 -1.496 312.36 0.054 3.571 44.82 0.271 -0.370 223.95 0.142 2.069 120.10
Ablated 0.401 0.000 306.72 0.056 0.000 45.28 0.270 0.000 221.81 0.145 0.000 121.46
SFM ($\alpha$=1.0) 0.361 9.858 277.56 0.053 4.171 44.08 0.225 16.924 186.23 0.132 9.100 111.14
SFM ($\alpha$=2.0) 0.299 25.541 229.43 0.048 13.306 39.16 0.187 30.931 153.08 0.115 20.484 96.02
SFM ($\alpha$=3.0) 0.249 37.931 191.49 0.043 22.749 34.88 0.157 41.895 129.02 0.100 30.753 84.20
SFM ($\alpha$=4.0) 0.207 48.486 156.78 0.037 34.246 29.74 0.131 51.576 107.90 0.086 40.559 72.56
SFM ($\alpha$=5.0) 0.157 60.825 120.07 0.027 50.968 22.14 0.110 59.266 90.74 0.076 47.605 63.74
@@
Baseline 0.972 -2.101 318.42 0.139 -6.107 49.00 0.681 1.304 243.40 0.365 1.617 133.92
Ablated 0.952 0.000 313.18 0.131 0.000 45.52 0.690 0.000 245.69 0.371 0.000 134.76
SFM ($\alpha$=1.0) 0.950 0.210 312.26 0.123 6.269 42.60 0.690 0.030 244.43 0.371 0.111 134.00
SFM ($\alpha$=2.0) 0.771 18.998 253.66 0.111 15.463 38.28 0.558 19.071 196.58 0.318 14.356 114.08
SFM ($\alpha$=3.0) 0.626 34.271 205.49 0.100 23.635 34.50 0.436 36.844 153.86 0.265 28.626 95.48
SFM ($\alpha$=4.0) 0.519 45.449 170.63 0.093 29.503 31.96 0.350 49.257 123.80 0.231 37.846 83.48
SFM ($\alpha$=5.0) 0.407 57.242 133.78 0.073 44.413 25.10 0.279 59.547 99.05 0.195 47.454 70.28
@@
Ablated 2.153 0.000 712.06 0.394 0.000 133.96 1.270 0.000 434.74 0.998 0.000 346.12
SFM ($\alpha$=1.0) 1.212 43.694 397.26 0.261 33.707 88.36 0.790 37.823 269.32 0.690 30.849 232.60
SFM ($\alpha$=2.0) 0.982 54.386 316.00 0.232 41.301 76.48 0.697 45.123 231.82 0.613 38.593 208.72
SFM ($\alpha$=3.0) 0.789 63.346 254.74 0.205 48.081 67.44 0.611 51.930 204.64 0.564 43.473 190.60
SFM ($\alpha$=4.0) 0.592 72.497 192.34 0.169 57.149 53.84 0.542 57.305 181.72 0.489 51.003 166.12
SFM ($\alpha$=5.0) 0.586 72.792 189.30 0.173 56.096 53.64 0.538 57.674 178.90 0.505 49.377 169.48
@@
Baseline 2.416 4.581 390.06 0.358 2.981 58.62 1.894 12.920 327.25 1.294 7.373 215.06
Ablated 2.532 0.000 395.25 0.369 0.000 58.91 2.175 0.000 346.90 1.397 0.000 223.37
SFM ($\alpha$=1.0) 1.745 31.093 275.92 0.283 23.449 45.68 1.454 33.143 232.67 1.061 24.069 170.57
SFM ($\alpha$=2.0) 1.426 43.667 218.17 0.250 32.196 39.63 1.211 44.326 192.97 0.917 34.395 146.30
SFM ($\alpha$=3.0) 1.110 56.182 173.28 0.219 40.682 34.85 1.001 53.952 158.57 0.787 43.679 126.53
SFM ($\alpha$=4.0) 0.963 61.967 149.95 0.186 49.642 29.98 0.887 59.217 142.15 0.720 48.469 116.09
SFM ($\alpha$=5.0) 0.960 62.096 149.90 0.185 49.875 29.96 0.888 59.190 141.79 0.727 47.990 116.15
@@
Ablated 1.170 0.000 834.46 0.188 0.000 137.56 0.699 0.000 514.30 0.539 0.000 403.36
SFM ($\alpha$=1.0) 0.621 46.934 421.90 0.125 33.704 86.84 0.396 43.324 277.72 0.337 37.591 234.70
SFM ($\alpha$=2.0) 0.503 56.980 335.96 0.108 42.475 75.28 0.336 51.894 237.88 0.286 46.910 207.94
SFM ($\alpha$=3.0) 0.391 66.558 267.38 0.092 51.292 63.82 0.295 57.811 207.58 0.257 52.252 184.24
SFM ($\alpha$=4.0) 0.300 74.320 205.89 0.073 61.121 51.72 0.259 62.958 182.17 0.229 57.595 163.48
SFM ($\alpha$=5.0) 0.302 74.212 205.67 0.074 60.825 51.74 0.258 63.061 182.23 0.229 57.523 163.06
Table 11: RTF results when using adaptive-step ODE solvers (model inference). $\overline{\text{RTF}}$ denotes the mean of RTF. Rate (%) denotes the relative speedup compared to the ablated model.
@@
Baseline 0.409 -1.489 0.055 3.509 0.273 -0.368 0.143 2.055
Ablated 0.403 0.000 0.057 0.000 0.272 0.000 0.146 0.000
SFM ($\alpha$=1.0) 0.363 9.809 0.055 4.017 0.226 16.814 0.133 8.976
SFM ($\alpha$=2.0) 0.300 25.430 0.050 12.866 0.188 30.735 0.117 20.230
SFM ($\alpha$=3.0) 0.251 37.773 0.045 22.063 0.159 41.639 0.102 30.399
SFM ($\alpha$=4.0) 0.208 48.282 0.038 33.261 0.133 51.268 0.088 40.104
SFM ($\alpha$=5.0) 0.159 60.578 0.029 49.541 0.112 58.914 0.077 47.076
@@
Baseline 0.975 -2.094 0.142 -5.185 0.684 1.441 0.369 1.600
Ablated 0.955 0.000 0.135 0.000 0.694 0.000 0.375 0.000
SFM ($\alpha$=1.0) 0.953 0.211 0.127 6.111 0.693 0.036 0.374 0.116
SFM ($\alpha$=2.0) 0.775 18.925 0.115 15.046 0.562 18.971 0.321 14.213
SFM ($\alpha$=3.0) 0.629 34.141 0.104 23.000 0.439 36.651 0.268 28.349
SFM ($\alpha$=4.0) 0.523 45.274 0.096 28.711 0.354 48.998 0.234 37.481
SFM ($\alpha$=5.0) 0.411 57.021 0.077 43.209 0.283 59.235 0.199 46.993
@@
Ablated 2.157 0.000 0.398 0.000 1.274 0.000 1.001 0.000
SFM ($\alpha$=1.0) 1.216 43.622 0.265 33.396 0.793 37.716 0.694 30.730
SFM ($\alpha$=2.0) 0.986 54.294 0.235 40.912 0.701 44.992 0.616 38.447
SFM ($\alpha$=3.0) 0.793 63.242 0.208 47.643 0.614 51.781 0.568 43.315
SFM ($\alpha$=4.0) 0.596 72.380 0.172 56.640 0.546 57.146 0.492 50.821
SFM ($\alpha$=5.0) 0.589 72.674 0.177 55.598 0.541 57.513 0.509 49.200
@@
Baseline 2.417 4.580 0.359 2.973 1.895 12.914 1.295 7.368
Ablated 2.533 0.000 0.370 0.000 2.176 0.000 1.398 0.000
SFM ($\alpha$=1.0) 1.746 31.076 0.284 23.353 1.455 33.124 1.062 24.035
SFM ($\alpha$=2.0) 1.428 43.644 0.252 32.070 1.212 44.301 0.918 34.360
SFM ($\alpha$=3.0) 1.111 56.155 0.220 40.536 1.003 53.922 0.788 43.638
SFM ($\alpha$=4.0) 0.964 61.937 0.187 49.471 0.888 59.184 0.721 48.424
SFM ($\alpha$=5.0) 0.961 62.066 0.186 49.705 0.889 59.157 0.728 47.946
@@
Ablated 1.171 0.000 0.190 0.000 0.700 0.000 0.541 0.000
SFM ($\alpha$=1.0) 0.622 46.870 0.126 33.406 0.397 43.221 0.338 37.472
SFM ($\alpha$=2.0) 0.505 56.903 0.110 42.113 0.338 51.779 0.288 46.775
SFM ($\alpha$=3.0) 0.393 66.472 0.093 50.872 0.296 57.683 0.259 52.102
SFM ($\alpha$=4.0) 0.302 74.225 0.075 60.633 0.260 62.820 0.230 57.432
SFM ($\alpha$=5.0) 0.303 74.116 0.075 60.336 0.260 62.924 0.230 57.360